# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v3

@@ II.2 Quantum geometric tensor
$X_{\mu}\coloneqq-\mathrm{i}\frac{\partial}{\partial\lambda^{\mu}}U(g(\bm{\lambda}))\biggl{|}_{\bm{\lambda}=\bm{0}}\quad$ (4)
@@
Note that in this case, the size of a QGT is determined only by the dimension of the Lie group, $\dim G$, independent of the dimensions of Hilbert spaces or the representations of the group. Since QGTs are Hermitian matrices, we can introduce a partial order $\mathcal{Q}^{\psi}\geq\mathcal{Q}^{\phi}$, which means that $\mathcal{Q}^{\psi}-\mathcal{Q}^{\phi}$ is positive semi-definite. We remark that this ordering is independent of the parametrization of the Lie group $G$ since both $\mathcal{Q}^{\psi}$ and $\mathcal{Q}^{\phi}$ transform as tensors under a coordinate transformation on the group.
@@ III Main results
[…]
$R(\psi\to\phi)=\sup\{r\geq 0\mid\forall i,\,\mathcal{Q}^{\mathcal{U}_{g_{i}}(\psi)}\geq r\mathcal{Q}^{\mathcal{U}_{g_{i}}^{\prime}(\phi)}\},$ (6)
where $g_{i}\in G_{i}$ is a representative element of each connected component $G_{i}$ of $G$. In particular, if $G$ is connected, then
$R(\psi\to\phi)=\sup\{r\geq 0\mid\mathcal{Q}^{\psi}\geq r\mathcal{Q}^{\phi}\}.$ (7)
@@
Since a compact Lie group has only a finite number of connected components (see, e.g., [64]), the conversion rate in Theorem 1 is expressed in terms of a finite number of matrix inequalities. In particular, Eq. (6) is independent of the choice of the elements $g_{i}$, and thus can be easily evaluated. Although Theorem 1 is of practical significance, the following formula is simple and theoretically useful, where all group elements are treated equally:
@@
The equivalence of Eq. (6) in Theorem 1 to Eq. (8) follows from the fact that if the matrix inequality holds at a point in a connected component of the group, then it also holds at any other point in the connected component. This is because the QGTs at different group elements in a connected component are interrelated by a congruence transformation independent of representation, as detailed in Appendix A.2. We note that the same argument for quantum Fisher information matrices instead of QGT can be found in [65].
@@
$R(\psi\to\phi)=2^{-\max_{i}D_{\max}\left(\mathcal{Q}^{\phi_{g_{i}}}\middle\|\mathcal{Q}^{\psi_{g_{i}}}\right)},$ (11)
where we have introduced the shorthand notations: $\psi_{g_{i}}\coloneqq\mathcal{U}_{g_{i}}(\psi)$ and $\phi_{g_{i}}\coloneqq\mathcal{U}_{g_{i}}^{\prime}(\phi)$. In particular, for a connected compact Lie group, this expression is further simplified to
@@
In the following, we prove Eq. (8) under the assumption that $U$ and $U^{\prime}$ are (non-projective) unitary representations of a compact Lie group $G$. The extension of the formula to any projective unitary representations is provided in Appendix A.3 by using the method in [59] that relates the conversion rate for projective unitary representations to (non-projective) unitary representations.
We remark that for any states $\rho$ and $\sigma$, Proposition 5 in [14] shows that $\mathrm{Sym}_{G}(\rho)\not\subset\mathrm{Sym}_{G}(\sigma)$ implies that the conversion from $\rho$ to $\sigma$ without error is impossible via a $G$-covariant channel. As its generalization to approximate asymptotic conversion, we prove Eq. (9) in Appendix A.4.
@@ III.1 U(1) group
In a study [24] on RTA for time-translation symmetry, the approximate asymptotic conversion rate among states having the same finite period was calculated, which generalizes prior studies [73, 74, 12]. When the Hamiltonians of the input and output systems are $H$ and $H^{\prime}$, the time-translation unitary operators are given by $e^{-\mathrm{i}Ht}$ and $e^{-\mathrm{i}H^{\prime}t}$. For pure states $\psi$ and $\phi$, the periods are defined as $\tau\coloneqq\inf\{t>0\mid e^{-\mathrm{i}Ht}\psi e^{\mathrm{i}Ht}=\psi\}$ and $\tau^{\prime}\coloneqq\inf\{t>0\mid e^{-\mathrm{i}H^{\prime}t}\phi e^{\mathrm{i}H^{\prime}t}=\phi\}$. In Theorem 1 in [24], it is proven that if $\tau=\tau^{\prime}$, the conversion rate is given by the ratio of the variances, i.e.,
$R(\psi\to\phi)=\frac{V(\psi,H)}{V(\phi,H^{\prime})}.$ (13)
We note that, when investigating the conversion among states with the same period, we can assume without loss of generality that the Hamiltonians have integer eigenvalues after appropriately redefining them [24, 75]. Therefore, this result corresponds to the case of $G=U(1)$.
@@ III.2 Reversible asymptotic conversion: Proof of the Marvian-Spekkens conjecture
For a connected compact Lie group $G$, pure states $\psi$ and $\phi$ are asymptotically reversibly convertible if and only if both of the following conditions (A) and (B) are satisfied:
@@
    There uniquely exists $r>0$ such that $\mathcal{Q}^{\psi}=r\mathcal{Q}^{\phi}$.
@@
Proof.
@@
Let us now review the statement of the Marvian-Spekkens conjecture and prove that it is equivalent to Corollary 2. Let $\bm{X}=\{X_{k}\}_{k}$ be the representation of a basis of the Lie algebra $\mathfrak{g}$. We define the symmetirized covariance matrix as
$\left(C_{\mathfrak{g}}(\psi)\right)_{kl}$ $\coloneqq\frac{1}{2}\braket{\psi|(X_{k}X_{l}+X_{l}X_{k})|\psi}$
$\quad-\braket{\psi|X_{k}|\psi}\braket{\psi|X_{l}|\psi}.$ (16)
The commutator subalgebra $\mathrm{i}[\mathfrak{g},\mathfrak{g}]$ is defined as the subalgebra spanned by $\mathrm{i}[L_{1},L_{2}]$ for all $L_{1},L_{2}\in\mathfrak{g}$. In general, the input and output Hilbert spaces $\mathcal{H}_{\mathrm{in}}$ and $\mathcal{H}_{\mathrm{out}}$ to which $\psi$ and $\phi$ belong are different. However, as shown in Appendix B in [76], by considering a larger Hilbert space $\mathcal{H}\coloneqq\mathcal{H}_{\mathrm{in}}\oplus\mathcal{H}_{\mathrm{out}}$ and a representation on it, it suffices to consider the case where the input and output Hilbert spaces are the same. The Marvian–Spekkens conjecture [16, 14] states that pure states $\psi$ and $\phi$ are asymptotically reversibly convertible in RTA for a connected compact Lie group $G$ if and only if all the following three conditions are satisfied:
@@
We remark that other results on reversible conversion rates [12, 38, 24] also follow since the Marvian–Spekkens conjecture has now been proven.
@@ III.4 Finite groups
Here, we analyze the approximate asymptotic conversion theory in RTA for a finite group. As proven in [59], the conversion rate diverges for a finite group. We show that this fact can also be derived from our formula in Eq. (6) by lifting a finite group to a compact Lie group by appending a trivial phase, though a finite group itself is not a Lie group  [^1] Formally, a finite group can be regarded as a Lie group since it can be viewed as a zero-dimensional smooth manifold. However, in our terminology, we consider Lie groups to have a dimension greater than zero..
@@
reproducing the result on the approximate asymptotic conversion rate in [59]. We remark that our construction of the conversion channels provides an intuitive explanation of the reason why the conversion rate for finite groups diverges, as will be explained at the end in Sec. V.3.
IV QGT as asymmetry monotone
In this section, we show that QGT is an asymmetry monotone both in one-shot and asymptotic conversions. As a consequence, we find that QGT provides a bound on the conversion rate.
IV.1 Monotonicity of QGT in one-shot conversion
We here prove the monotonicity of the QGT under conversion without error using a $G$-covariant channel. For this purpose, let us first relate the QGT to monotone metrics [78, 79], which are Riemannian metrics on state space contracting under information processing. The pioneering research on monotone metrics initiated by Morozova and Chentsov [78] was later completed by Petz [79], proving that monotone metrics are in a one-to-one correspondence with operator monotone functions. Here, a function $f:[0,\infty)\to[0,\infty)$ is called an operator monotone iff $0\leq A\leq B$ implies $f(A)\leq f(B)$. Petz’s monotone metric associated with an operator monotone function $f$ is defined by
$\braket{A,B}_{f,\rho}\coloneqq\mathrm{Tr}\left(A^{\dagger}m_{f}^{-1}(\mathcal{R}_{\rho},\mathcal{L}_{\rho})(B)\right),$ (22)
where $m_{f}(x,y)\coloneqq yf(x/y)$, $\mathcal{R}_{\rho}(A)\coloneqq A\rho$ and $\mathcal{L}_{\rho}(A)\coloneqq\rho A$. Here, when $\rho$ is not invertible, following the argument in [80], $m_{f}^{-1}(\mathcal{R}_{\rho},\mathcal{L}_{\rho})$ denotes the Moore-Penrose inverse, such that
$|,\displaystyle m_{f}^{-1}(\mathcal{R}_{\rho},\mathcal{L}_{\rho})(B)\coloneqq\sum_{k,l;m_{f}(p_{k},p_{l})>0}\frac{\braket{l|B|k}}{m_{f}(p_{k},p_{l})}\ket{l}\bra{k}$, (23)
where we used the eigenvalue decomposition of a state $|\rho=\sum_{k}p_{k}\ket{k}\bra{k}$. For a linear operator $A$, we denote $\|A\|_{f,\rho}\coloneqq\sqrt{\braket{A,A}_{f,\rho}}$.
An additional symmetry condition $f(t)=tf(t^{-1})$ is often imposed to the monotone function, which is equivalent to $\braket{A,B}_{f,\rho}=\braket{B^{\dagger},A^{\dagger}}_{f,\rho}$ for any linear operators $A$ and $B$. A famous example is $f_{\mathrm{SLD}}(x)\coloneqq(1+x)/2$ associated with the symmetric logarithmic derivative (SLD) metric, which plays a central role in quantum estimation theory. However, such a symmetric monotone metric does not contain the information of the anti-symmetric part of the QGT. Therefore, we do not impose the symmetry condition.
In order to investigate the properties of QGTs, we consider a family of operator monotone functions given by
@@ IV.1 Monotonicity of QGT in one-shot conversion
Note that this family contains $f_{\mathrm{SLD}}$ as a special case for $q=1/2$. Let us show that the QGT can be obtained as a limit of Petz’s monotone metric for $f_{q}$. Since $m_{f_{q}}(p_{k},p_{l})=(1-q)p_{l}+qp_{k}$, for any pure state $\psi$ and any linear operator $O$, it holds
@@
where we have defined a generalized variance $V$ by
$V(\psi,O)\coloneqq\braket{\psi|O(I-\psi)O^{\dagger}|\psi}$ (26)
for a pure state $\psi$ and a linear operator $O$. Here, following a standard convention, we described a pure state and a unit vector by the same Greek letter, i.e., $|\psi=\ket{\psi}\bra{\psi}$. Note that when $O$ is Hermitian, $V$ is equivalent to the ordinary variance.
When $O$ is given by $O\coloneqq\bm{\gamma}^{\dagger}\bm{X}\coloneqq\sum_{\mu=1}^{\dim G}\gamma^{*\mu}X_{\mu}$ for $\bm{\gamma}\in\mathbb{C}^{\dim G}$ and $\bm{X}=\{X_{\mu}\}_{\mu=1}^{\dim G}$ defined in Eq. (4), we get
$\|\mathrm{i}[\psi,O]\|_{f_{q},\psi}^{2}=\bm{\gamma}^{\dagger}\left(\frac{1}{1-q}\mathcal{Q}^{\psi}+\frac{1}{q}\left(\mathcal{Q}^{\psi}\right)^{*}\right)\bm{\gamma}.$ (27)
Therefore, we find
$\lim_{q\to 1^{-}}f_{q}(0)\|\mathrm{i}[\psi,O]\|_{f_{q},\psi}^{2}=\bm{\gamma}^{\dagger}\mathcal{Q}^{\psi}\bm{\gamma},$ (28)
where we have used $f_{q}(0)=1-q>0$.
We stress again that the operator monotone function satisfying the symmetry condition is not suitable for investigating the monotonicity of QGT. Indeed, for operator monotone functions $f^{(\mathrm{s})}$ satisfying $f^{(\mathrm{s})}(0)>0$ and the symmetry condition $f^{(\mathrm{s})}(t)=tf^{(\mathrm{s})}(t^{-1})$, from the definition in Eq. (22), we find
$\frac{f^{(\mathrm{s})}(0)}{2}\|\mathrm{i}[\psi,O]\|_{f^{(\mathrm{s})},\psi}^{2}=\bm{\gamma}^{\dagger}\left(\frac{1}{2}\left(\mathcal{Q}^{\psi}+\left(\mathcal{Q}^{\psi}\right)^{*}\right)\right)\bm{\gamma},$ (29)
for $O\coloneqq\bm{\gamma}^{\dagger}\bm{X}$, meaning that it captures only the symmetric part of $\mathcal{Q}^{\psi}$.
This equation can be viewed as a special case of Eq. (29), where $\bm{\gamma}$ is a real vector.
The metric $\braket{\cdot,\cdot}_{f_{q},\rho}$ is monotonic under a quantum channel in the following sense:
@@
In literature, the proof of the monotonicity of monotone metric is sometimes provided under the assumption that $\rho$ is invertible. Lemma 3 is a special case of Theorem 6.1 in [82], which is valid not only for invertible states but also for non-invertible states. The details of the proof of Lemma 3 and its relation to Theorem 6.1 in [82] are presented in Appendix B.1.
The monotonicity of the QGT in RTA is proven by using Lemma 3. Suppose that a state $\rho$ is convertible without error to another state $\sigma$ via a $G$-covariant channel $\mathcal{E}$, i.e., $\mathcal{E}(\rho)=\sigma$. Since the $G$-covariance of the channel implies $\mathcal{E}(\mathcal{U}_{g}(\rho))=\mathcal{U}_{g}^{\prime}(\sigma)$, its derivative with respect to the parameter $\bm{\lambda}$ for $g\in G$ yields $\mathcal{E}\left(\mathrm{i}[\rho,X_{\mu}]\right)=\mathrm{i}[\sigma,X_{\mu}^{\prime}]$, where $X_{\mu}$ is given in Eq. (4), and $X_{\mu}^{\prime}$ is defined similarly by differentiating the projective unitary representation $U^{\prime}$ of $G$ on the output system instead of $U$.
By using the linearity of $\mathcal{E}$, we get $\mathcal{E}(\mathrm{i}[\rho,O])=\mathrm{i}[\mathcal{E}(\rho),O^{\prime}]$ for $O\coloneqq\bm{\gamma}^{\dagger}\bm{X}$ and $O^{\prime}\coloneqq\bm{\gamma}^{\dagger}\bm{X}^{\prime}$ for any $\bm{\gamma}\in\mathbb{C}^{\dim G}$. The monotonicity in Lemma 3 implies
@@
Let us now apply this inequality to the case where $\rho$ and $\sigma$ are pure states, which we denote by $\psi$ and $\phi$, respectively. Multiplying Eq. (32) by $f_{q}(0)=1-q>0$ and taking the limit of $q\to 1^{-}$, Eq. (28) implies, $\bm{\gamma}^{\dagger}\mathcal{Q}^{\psi}\bm{\gamma}\geq\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}$ for any $\bm{\gamma}\in\mathbb{C}^{\dim G}$, or equivalently,
@@
IV.2 Asymptotic discontinuity of QGT
When the projective unitary representation of $G$ on a single system is given by $U$, the representation of $G$ on $N$ i.i.d. copies of the system is given by $U^{\otimes N}$. Therefore, for an i.i.d. pure state $\psi^{\otimes N}$, the QGT is additive, i.e., $\mathcal{Q}^{\psi^{\otimes N}}=N\mathcal{Q}^{\psi}$ for any positive integer $N$. When $\{\psi^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}$, combined with the monotonicity of QGT in Eq. (33), one would expect that the asymptotic rates of the QGTs for these sequences, i.e., $\lim_{N\to\infty}\frac{1}{N}\mathcal{Q}^{\psi^{\otimes N}}$ and $\lim_{N\to\infty}\frac{1}{N}\mathcal{Q}^{\phi^{\otimes\lfloor rN\rfloor}}$, are monotonically decreasing, implying that

### v4

@@ II.2.1 Definition and basic properties
We remark that the QGT transforms as a $(0,2)$-tensor under smooth reparametrizations of the parameters $\lambda^{\mu}$. The indices $\mu$ and $\nu$ label tensor components: a superscript index indicates a contravariant component, whereas a subscript index indicates a covariant component under such reparametrizations. In this paper, we use tensor notation to clarify the transformation properties, even though we do not explicitly change coordinates. Hereafter, we employ the Einstein summation convention: whenever an index appears once as an upper index and once as a lower index, summation over that index is implied.
@@
Since the QGT is Hermitian, i.e., $\mathcal{Q}_{\mu\nu}^{\psi,\xi}=(\mathcal{Q}_{\nu\mu}^{\psi,\xi})^{*}$, its real and imaginary parts define symmetric and antisymmetric tensors, respectively. Specifically, the real part $\mathrm{Re}(\mathcal{Q}_{\mu\nu}^{\psi,\xi})$ is symmetric and corresponds to the quantum Fisher information matrix, while twice the imaginary part, $2\,\mathrm{Im}(\mathcal{Q}_{\mu\nu}^{\psi,\xi})$, is anti-symmetric and corresponds to the Berry curvature (for further details, see, e.g., [44, 45, 78, 79]).
II.2.2 QGT in the RTA
[…]
$X_{\mu}\coloneqq-\mathrm{i}\frac{\partial}{\partial\lambda^{\mu}}U(g(\lambda))\biggl|_{\lambda=0}\quad$ (6)
for $\mu=1,\cdots,\dim G$. In other words, by introducing the derived representation $L$ of the Lie algebra associated with $U$ by $L(A)\coloneqq-\mathrm{i}\left.\frac{\mathrm{d}}{\mathrm{d}t}U(e^{\mathrm{i}tA})\right|_{t=0}$, the operator $X_{\mu}$ is equal to $L(A_{\mu})$.
@@ II.2.2 QGT in the RTA
Note that in this case, the size of the QGT is determined only by the dimension of the Lie group, $\dim G$, independent of the dimensions of Hilbert spaces or the representations of the group. Since QGTs are Hermitian matrices, we can introduce a partial order $\mathcal{Q}^{\psi}\geq\mathcal{Q}^{\phi}$, which means that $\mathcal{Q}^{\psi}-\mathcal{Q}^{\phi}$ is positive semi-definite. We remark that this ordering is independent of the parametrization of the Lie group $G$ since both $\mathcal{Q}^{\psi}$ and $\mathcal{Q}^{\phi}$ transform as tensors under a coordinate transformation on the group.
[…]
@@ III Main results
Let $U$ and $U^{\prime}$ be projective unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. The conversion rate from a pure state $\psi\in\mathcal{P}(\mathcal{H})$ to another pure state $\phi\in\mathcal{P}(\mathcal{H}^{\prime})$ is given by
$R(\psi\to\phi)=\sup\{r\geq 0\mid\mathcal{Q}^{\psi}\geq r\mathcal{Q}^{\phi}\}$ (8)
if $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$, and $R(\psi\to\phi)=0$ otherwise.
@@
Although Theorem 1 is of practical significance, the following formula is also theoretically useful, where all group elements are treated equally:
@@
The equivalence of Eq. (8) in Theorem 1 to Eq. (9) follows from the fact that if the matrix inequality holds at a point in a compact Lie group, then it also holds at any other point in the group. This is because the QGTs at different group elements are interrelated by a congruence transformation independent of representation, as detailed in Appendix B.2. We note that the same argument for quantum Fisher information matrices instead of QGT can be found in [80].
As we show in Sec. IV.1, the QGT is an asymmetry measure: it is non-increasing under any $G$-covariant channel with respect to the partial order defined by matrix inequalities, i.e., the Löwner order. Equation (8), therefore, identifies the QGT as the complete measure of symmetry breaking.
This result carries an important conceptual consequence. Resource measures are typically formulated as real-valued scalar functions, while operator-valued resource measures have recently been considered in [81, 80]. Our formula shows that an operator, namely the QGT matrix, governs the asymptotic conversion problem in the RTA. This establishes that the operator-valued formulation is not merely a generalization of the scalar formulation, but a necessary extension for resolving this central problem.
Within this operator-valued paradigm, scalar-valued measures are recovered as the special case of $1\times 1$ matrices. As we shall see explicitly, this reduction occurs, for example, for $U(1)$ symmetry.
@@
Note that the optimization in Eq. (8) can be written explicitly as
$\begin{aligned} \text{maximize}\quad&r\\ \text{subject to}\quad&\mathcal{Q}^{\psi}-r\,\mathcal{Q}^{\phi}\geq 0,\\ &r\geq 0.\end{aligned}$ (13)
Since the constraint $\mathcal{Q}^{\psi}-r\,\mathcal{Q}^{\phi}\geq 0$ is a linear matrix inequality in the scalar variable $r$, this problem is a semidefinite program. Moreover, by using the Moore–Penrose inverse $\mathcal{Q}^{\psi+}$ of the QGT $\mathcal{Q}^{\psi}$, we obtain the following closed-form expression:
$R(\psi\to\phi)$
$\quad=\begin{cases}+\infty&(\text{if }\mathcal{Q}^{\phi}=0)\\ 0&(\text{if }\ker(\mathcal{Q}^{\psi})\not\subset\ker(\mathcal{Q}^{\phi}))\\ \displaystyle\frac{1}{\bigl\|(\mathcal{Q}^{\psi+})^{1/2}\mathcal{Q}^{\phi}(\mathcal{Q}^{\psi+})^{1/2}\bigr\|_{\infty}}&(\text{otherwise})\end{cases}$ (14)
where $\|A\|_{\infty}$ denotes the operator norm of $A$.
Key ideas for proving Eq. (9), under the assumption that $U$ and $U^{\prime}$ are (non-projective) unitary representations of a compact Lie group $G$, are explained in Sec. IV, while a rigorous proof is provided in Appendices C and D. The extension of the formula to any projective unitary representations is provided in Appendix B.3 by using the method in [64] that relates the conversion rate for projective unitary representations to (non-projective) unitary representations.
We remark that for any states $\rho$ and $\sigma$, Proposition 5 of [14] shows that if $\mathrm{Sym}_{G}(\rho)\not\subset\mathrm{Sym}_{G}(\sigma)$, then an exact conversion from $\rho$ to $\sigma$ via a $G$-covariant channel is impossible. As a generalization of this observation to the asymptotic setting with vanishing error, we prove Eq. (10) in Appendix B.4. In Appendix K, we propose a way to circumvent this limitation by using additional resource states.
@@ III.1 U(1) group
In a study [24] on RTA for time-translation symmetry, the asymptotic conversion rate among states having the same finite period was calculated, which generalizes prior studies [89, 90, 12]. When the Hamiltonians of the input and output systems are $H$ and $H^{\prime}$, the time-translation unitary operators are given by $e^{-\mathrm{i}Ht}$ and $e^{-\mathrm{i}H^{\prime}t}$. For pure states $\psi$ and $\phi$, the periods are defined as $\tau\coloneqq\inf\{t>0\mid e^{-\mathrm{i}Ht}\psi e^{\mathrm{i}Ht}=\psi\}$ and $\tau^{\prime}\coloneqq\inf\{t>0\mid e^{-\mathrm{i}H^{\prime}t}\phi e^{\mathrm{i}H^{\prime}t}=\phi\}$. In Theorem 1 in [24], it is proven that if $\tau=\tau^{\prime}$, the conversion rate is given by the ratio of the variances, i.e.,
$R(\psi\to\phi)=\frac{V(\psi,H)}{V(\phi,H^{\prime})},$ (15)
where we defined $V(\psi,H)\coloneqq\braket{\psi|H^{2}|\psi}-\braket{\psi|H|\psi}^{2}$. We note that, when investigating the conversion among states with the same period, we can assume without loss of generality that the Hamiltonians have integer eigenvalues after appropriately redefining them [24, 91]. Therefore, this result corresponds to the case of $G=U(1)$.
@@ III.2 Reversible asymptotic conversion: Proof of the Marvian-Spekkens conjecture
For a compact Lie group $G$, pure states $\psi$ and $\phi$ are asymptotically reversibly convertible if and only if both of the following conditions (A) and (B) are satisfied:
@@
    There exists a unique $r>0$ such that $\mathcal{Q}^{\psi}=r\mathcal{Q}^{\phi}$.
@@
Here, in condition (B), the requirement that $r$ be unique excludes the case where $\mathcal{Q}^{\psi}=\mathcal{Q}^{\phi}=0$, for which the proportionality constant is not unique. In this case, one has $R(\psi\to\phi)=R(\phi\to\psi)=\infty$ since $\psi$ and $\phi$ are symmetric, and hence these states formally fail to satisfy the reversibility condition $R(\psi\to\phi)R(\phi\to\psi)=1$.
Proof of Corollary 2.
@@
Let us now review the statement of the Marvian-Spekkens conjecture and prove that it is equivalent to Corollary 2. Let $\{X_{\mu}\}_{\mu}$ be the derived representation of a basis of the Lie algebra $\mathfrak{g}$. We define the symmetirized covariance matrix as
$\left(C_{\mathfrak{g}}(\psi)\right)_{\mu\nu}$ $\coloneqq\frac{1}{2}\braket{\psi|(X_{\mu}X_{\nu}+X_{\nu}X_{\mu})|\psi}$
$\quad-\braket{\psi|X_{\mu}|\psi}\braket{\psi|X_{\nu}|\psi}.$ (18)
The commutator subalgebra $\mathrm{i}[\mathfrak{g},\mathfrak{g}]$ is defined as the subalgebra spanned by $\mathrm{i}[L_{1},L_{2}]$ for all $L_{1},L_{2}\in\mathfrak{g}$. In general, the input and output Hilbert spaces $\mathcal{H}_{\mathrm{in}}$ and $\mathcal{H}_{\mathrm{out}}$, to which $\psi$ and $\phi$ belong respectively, are different. However, as shown in Appendix B in [92], by considering a larger Hilbert space $\mathcal{H}\coloneqq\mathcal{H}_{\mathrm{in}}\oplus\mathcal{H}_{\mathrm{out}}$ and a representation on it, it suffices to consider the case where the input and output Hilbert spaces are the same. The Marvian–Spekkens conjecture [16, 14] states that pure states $\psi$ and $\phi$ are asymptotically reversibly convertible in RTA for a connected compact Lie group $G$ if and only if all the following three conditions are satisfied:
@@
We remark that other results on reversible conversion rates [12, 38, 24] also follow since the Marvian–Spekkens conjecture has now been proven. It is also worth emphasizing that, while the Marvian–Spekkens conjecture was originally formulated for connected compact Lie groups, Corollary 2 holds for non-connected compact Lie groups as well.
@@ III.4 Finite groups
Here, we analyze the asymptotic conversion theory in RTA for a finite group. As proven in [64], the conversion rate diverges for a finite group. We show that this fact can also be derived from our formula in Eq. (8) by lifting a finite group to a compact Lie group by appending a trivial phase, though a finite group itself is not a Lie group  [^1] Formally, a finite group can be regarded as a Lie group since it can be viewed as a zero-dimensional smooth manifold. However, in our terminology, we consider Lie groups to have a dimension greater than zero..
@@
reproducing the result on the asymptotic conversion rate in [64]. We remark that our construction of the conversion channels provides an intuitive explanation of the reason why the conversion rate for finite groups diverges, as will be explained at the end in Sec. D.1.2.
IV Proof sketch of Theorem 1
In this section, we provide a proof sketch of Theorem 1, and explain the key ideas. A fully detailed, rigorous proof can be found in Appendices C and D.
IV.1 QGT as asymmetry monotone
A key property of the QGT is its monotonicity as an asymmetry measure, in the form of a matrix inequality. We remark that resource monotones in quantum resource theories are usually given by real-valued functions on states that do not increase under free operations, thereby providing necessary conditions for convertibility. Recent studies [81, 80] have extended this viewpoint to matrix-valued monotones, which yield necessary conditions for convertibility formulated via partial orders induced by matrix inequalities. Our study aligns with these works in that the QGT is an asymmetry monotone expressed through a matrix inequality.
To prove the monotonicity of the QGT, we relate it to a family of Petz’s monotone metrics [65, 66], which are Riemannian metrics on the state space that contract under information processing. To introduce Petz’s monotone metrics, we begin with the notion of operator monotone functions. A function $f:[0,\infty)\to[0,\infty)$ is called operator monotone if and only if $0\leq A\leq B$ implies $f(A)\leq f(B)$. Following [94], Petz’s monotone metric associated with $f$ is defined by
$\braket{A,B}_{f,\rho}\coloneqq\sum_{k,l:\,m_{f}(p_{k},p_{l})>0}\frac{\braket{k|A^{\dagger}|l}\braket{l|B|k}}{m_{f}(p_{k},p_{l})},$ (24)
where $m_{f}(x,y)\coloneqq yf(x/y)$ and $|\rho=\sum_{k}p_{k}\ket{k}\bra{k}$ is the eigenvalue decomposition. For a linear operator $A$, we define the corresponding norm by $\|A\|_{f,\rho}\coloneqq\sqrt{\braket{A,A}_{f,\rho}}$.
We remark that an additional symmetry condition $f(t)=tf(t^{-1})$ is often imposed on operator monotone functions [66, 95]. However, as discussed in Appendix C.1, the monotone metric associated with a symmetric operator monotone function does not capture the anti-symmetric part of the QGT. We therefore do not impose this symmetry condition.
A central object in our study is a one-parameter family of operator monotone functions $f_{q}$, defined by
@@ IV.1 QGT as asymmetry monotone
Since $m_{f_{q}}(p_{k},p_{l})=(1-q)p_{l}+qp_{k}$, for any pure state $\psi$ and any linear operator $O$, we obtain
@@
where we have defined the generalized variance $V$ by $V(\psi,O)\coloneqq\braket{\psi|O(I-\psi)O^{\dagger}|\psi}$. Note that when $O$ is Hermitian, $V$ coincides with the ordinary variance.
To relate the QGT to the above norm, let $O=\gamma^{\dagger}X\coloneqq\gamma^{\mu*}X_{\mu}$, where $\{X_{\mu}\}_{\mu=1}^{\dim G}$ are generators defined in Eq. (6) and $\gamma\in\mathbb{C}^{\dim G}$. From Eq. (26), we obtain
$\|\mathrm{i}[\psi,O]\|_{f_{q},\psi}^{2}=\gamma^{\dagger}\left(\frac{1}{1-q}\mathcal{Q}^{\psi}+\frac{1}{q}\left(\mathcal{Q}^{\psi}\right)^{*}\right)\gamma,$ (27)
which implies
$\lim_{q\to 1^{-}}f_{q}(0)\|\mathrm{i}[\psi,O]\|_{f_{q},\psi}^{2}=\gamma^{\dagger}\mathcal{Q}^{\psi}\gamma.$ (28)
As detailed in Appendix C.2.1, the norm $\|\cdot\|_{f,\rho}$ is monotonic under a quantum channel in the following sense:
@@
The monotonicity of the QGT in RTA is proven by using Lemma 3. Suppose that a state $\rho$ is convertible without error to another state $\sigma$ via a $G$-covariant channel $\mathcal{E}$. Since the $G$-covariance of the channel implies $\mathcal{E}(\mathcal{U}_{g}(\rho))=\mathcal{U}_{g}^{\prime}(\sigma)$, its derivative with respect to the parameter $\lambda\in\mathbb{R}^{\dim G}$ for $g\in G$ yields $\mathcal{E}\left(\mathrm{i}[\rho,X_{\mu}]\right)=\mathrm{i}[\sigma,X_{\mu}^{\prime}]$, where $X_{\mu}$ is given in Eq. (6), and $X_{\mu}^{\prime}$ is defined similarly by differentiating the projective unitary representation $U^{\prime}$ of $G$ on the output system instead of $U$.
By using the linearity of $\mathcal{E}$, we get $\mathcal{E}(\mathrm{i}[\rho,O])=\mathrm{i}[\mathcal{E}(\rho),O^{\prime}]$ for $O\coloneqq\gamma^{{\dagger}}X$ and $O^{\prime}\coloneqq\gamma^{\dagger}X^{\prime}$ for any $\gamma\in\mathbb{C}^{\dim G}$. The monotonicity in Lemma 3 implies
@@
Let us now apply this inequality to the case where $\rho$ and $\sigma$ are pure states, which we denote by $\psi$ and $\phi$, respectively. Multiplying Eq. (30) by $f_{q}(0)=1-q>0$ and taking the limit of $q\to 1^{-}$, Eq. (28) implies, $\gamma^{\dagger}\mathcal{Q}^{\psi}\gamma\geq\gamma^{\dagger}\mathcal{Q}^{\phi}\gamma$ for any $\gamma\in\mathbb{C}^{\dim G}$, or equivalently,
@@
We now turn to the i.i.d. case. For $N$ i.i.d. copies of a system, the QGT is additive: $\mathcal{Q}^{\psi^{\otimes N}}=N\mathcal{Q}^{\psi}$. When $\{\psi^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}$, monotonicity suggests that the asymptotic QGT rates decrease, yielding